OpenStudy (anonymous):

Fine the x and y intercepts of -2x +3y +15

OpenStudy (anonymous):

First put the equation into slope-intercept form y=mx+b

OpenStudy (anonymous):

3y=-2x+15 then divide by 3. y=-2/3x+5

OpenStudy (anonymous):

b is the y-int so 5 is the y int

OpenStudy (anonymous):

to find the x intercept, just set y equal to zero and solve the equation

OpenStudy (jonnyvonny):

X- intercepts are roots, which is when a point touches the x-axis (the horizontal axis). This happens when, after you put the x-value in, the outcome is 0. Y- intercepts occur when a point touches the y-axis (the vertical line). This happens when you PUT 0 as the x-value; the answer you get is the y-intercept. So first we must get y alone. We do this by following reverse PEMDAS; we subtract and add first then divide ect. Assuming that the function is equal to 0, we first subtract "15" and add "2x", then divide the BOTH values by 3. \[-2x +3y +15=0\]\[3y= 2x-15\]\[y=2x/3-15/3\] The position of "2x" and "-15" are IRREVELANT, just make sure that the negative is in the proper value. We may simplify "15/3" to "5". \[y= 2x/3-5\] Lets find the y-intercept first because its the easiest, imo. To do this, we set the x-value to 0 and simplify. \[y=(2*0/3) -5\]\[y=-5\] This means when the x-value is 0, the y-value is -5; the point (0,-5). Now the x-intercept. The x-intercept occurs when the y-value is 0, hence, the point intercepts, or touches, the x-axis. Now we set the y-value to 0, and solve for x; get x alone. Once again, we follow reverse PEMDAS for this, add 5 over, multiply by 3, then divide by 2. \[0= 2x/3-5 \]\[5=2x/3\]\[15=2x \]\[15/2=x\] Thus, the x-intercept is (15/2,0). And the y-intercept is (-5,0). I hope this helped, feel free to pst me if you need more assistance. :)